# Gravitational Waves on Kerr Black Holes I: Reconstruction of Linearized Metric Perturbations

### v1

@@ 1 Introduction and summary
As an aside, we briefly comment on the origin of the two “miracles” uncovered by Teukolsky: 1) the Kerr metric perturbations $h_{\mu\nu}$ are encoded in two gauge-invariant Weyl scalars (1.4), each obeying the same decoupled scalar equation (1), and 2) this equation is completely separable. In fact, Teukolsky showed that the first “miracle” occurs in any algebraically special spacetime of Petrov type D [7]: in this sense, therefore, the reduction of the spin-$s$ field equations to a pair of decoupled scalar equations on $\Psi^{(s)}$ can be viewed as a consequence of algebraic specialness. [^2] In general, this decoupling property only holds for fixed $|s|$. When perturbations of multiple spins are present, their equations may not fully decouple: for example, the Kerr-Newman spacetime is of Petrov type D, but the Teukolsky equations for electromagnetic ($|s|=1$) and gravitational ($|s|=2$) perturbations do not decouple. In Schwarzschild, the second “miracle”—the separability of the spin-$s$ field equations—follows from spherical symmetry. This symmetry is broken in Kerr, but as Carter discovered [8], it is replaced by a “hidden” symmetry in the form of an irreducible rank-2 Killing tensor $K_{\mu\nu}$, whose existence implies the separability of the scalar wave equation. Indeed, the Killing equation for $K_{\mu\nu}$ implies that the differential operator $\mathcal{K}=\nabla^{\mu}K_{\mu\nu}\nabla^{\nu}$ commutes with the Laplacian $\nabla^{2}=\nabla^{\mu}g_{\mu\nu}\nabla^{\nu}$, and both commute with the action of the Killing vectors $\mathop{}\!\partial_{t}$ and $\mathop{}\!\partial_{\phi}$ generating the isometries of Kerr (stationarity and axisymmetry, respectively); since these four differential operators all commute, they are simultaneously diagonalizable, and their joint eigenvectors are precisely the separated eigenmodes $\Psi_{\omega\ell m}^{(0)}=e^{-i\omega t+im\phi}R_{\omega\ell m}^{(0)}(r)S_{\omega\ell m}^{(0)}(\theta)$; [^3] The eigenvalues corresponding to $i\mathop{}\!\partial_{t}$, $i\mathop{}\!\partial_{\phi}$, $\nabla^{2}$, and $\mathcal{K}$ are respectively the energy $\omega$, azimuthal angular momentum $m$, invariant mass $\mu^{2}$, and separation constant $\lambda_{\omega\ell m}^{(0)}$ of the mode solution to $\mathopen{}\mathclose{{}\left(\nabla^{2}-\mu^{2}}\right)\Psi_{\omega\ell m}^{(0)}=0$. Likewise, the geodesic equation (which is the eikonal limit of the wave equation) is completely integrable due to the existence of four conserved quantities associated with these operators (e.g., $\mathcal{K}_{\mu\nu}p^{\mu}p^{\nu}$ is the Carter constant). see, e.g.,[9] for a more detailed review. More generally, the separability of the spin-$s$ field equations can be derived from the existence of commuting differential operators, all of which are “hidden” in the sense that they are associated with symmetries of higher rank, rather than being geometrically realized as isometries. [^4] The symmetry operators that separate the spin-$s$ equations in Kerr are derived by Carter and McLenaghan [10] for $|s|=\frac{1}{2}$ and by Torres del Castillo [11] for $|s|=1$. Aksteiner and Bäckdahl [12] tackled the case $|s|=2$. While satisfying, symmetry-based arguments do not fully explain these “miracles”: for one, the existence of “hidden” symmetries is sufficient but not necessary for separability. [^5] The C-metric describing a pair of accelerating black holes [13] provides a counter-example: it is algebraically special of Petrov type D and therefore admits a conformal Killing tensor, but it does not possess an exact one [14]. We are not aware of any method to derive separability from conformal symmetries and their differential operators. Dudley and Finley [15] proved that all spacetimes in the Plebański-Demański class [16], which includes all vacuum type D solutions, share these two properties (decoupling and separability).
@@
Although knowledge of the Weyl scalars $\psi_{0}$ and/or $\psi_{4}$ is adequate for many purposes (such as the computation of energy and angular momentum fluxes), the full metric perturbation $h_{\mu\nu}$ is often required to answer other questions about the perturbed spacetime geometry (such as the shear of the perturbed horizon), or to push perturbation theory beyond first order [18, 19]. It is therefore useful to reconstruct the specific metric perturbation $h_{\mu\nu}$ associated with a given Weyl scalar. Various methods for carrying out this metric reconstruction have been developed, but the explicit form of the metric perturbation $h_{\mu\nu}$ that corresponds to a single mode of $\psi_{0}$ or $\psi_{4}$ has—to our knowledge—never been published. The main purpose of this paper is to close this gap in the literature. Before describing our results, we briefly summarize preexisting work.
@@
$\text{ORG}:\qquad n^{\mu}h_{\mu\nu}^{\rm ORG}=0,\qquad g^{\mu\nu}h_{\mu\nu}^{\rm IRG}=0.$ (1.21)
@@ 1.1 Mode inversion and comparison with previous work
To carry out the metric reconstruction procedure explicitly, one must invert the relations (1.22) and (1.23) to obtain the Hertz potentials $\Psi_{\rm H}$ and $\Psi_{\rm H}^{\prime}$ in terms of the Weyl scalars $\psi_{0}$ and $\psi_{4}$. Although this inversion is difficult—or perhaps even impossible—to perform in full generality, it is straightforward (if messy) to invert these equations at the mode level, that is, for the case where $\Psi^{(+2)}=\psi_{0}$ or $\Psi^{(-2)}=\zeta^{4}\psi_{4}$ consists of a single mode (1.9b). For real frequencies $\omega$, the results of this mode inversion have previously appeared in the literature, though scattered across multiple papers, as we review below. In this paper, we tackle the general case of arbitrary (and possibly complex) frequency $\omega$, and completely invert all of the relations (1.22) and (1.23).
@@
In their excellent review, Pound and Wardell [28] inverted precisely the other two relations. Their Eq. (106) inverts our Eq. (1.23a) for $\Psi_{\rm H}^{\prime}$ (in their notation, $\psi^{\rm ORG}$) in terms of $\Psi^{(+2)}=\psi_{0}$, which they denote $\psi_{2}$—see their Eq. (81). Likewise, their Eq. (107) inverts our Eq. (1.22b) for $\Psi_{\rm H}$ (in their notation, $\psi^{\rm IRG}$) in terms of $\Psi^{(-2)}=\zeta^{4}\psi_{4}$, which they label $\psi_{-2}$—see their Eq. (82). Together, these three papers [26, 27, 28] completely invert all four of the relations (1.22) and (1.23), but only for the case of real frequency $\omega$, which can be a serious limitation for many applications.
Here, we invert all of these relations without assuming that the frequency is real, so that our results also hold for complex $\omega$. As a final technical point, although we treat all four equations, there is another redundancy inherent in mode inversion that can simplify the problem. The key observation is that, even if one wishes to use both $\psi_{0}$ and $\psi_{4}$, and work in both IRG and ORG, it is still enough to invert only two of the relations (1.22) and (1.23): that is, it suffices to invert Eqs. (1.22b) and (1.23a) only, as Pound and Wardell did, or else Eqs. (1.22a) and (1.23b) only.
@@
As described above, the second step was worked out for real frequency $\omega$ by multiple authors [26, 27, 28]. Here, our main contribution is to extend the mode inversion to complex frequency $\omega$. We also derive direct relations between modes of $\psi_{0}$ and $\psi_{4}$, which are to our knowledge novel.
@@ 1.2 Structure of the paper
Lastly, we summarize the theory of the confluent Heun function, and carefully derive the quantization conditions for its eigenvalues, in App. E. We also discuss the orthogonality of the confluent Heun functions and derive an analytic formula for their normalization, following the work of Becker [33] on the full Heun equation. Finally, we apply these results to the angular ODE (1.11) in App. F and to the radial ODE (1.10) in App. G, which collect more technicalities.
@@ 1.3 Present limitations and future directions
While we strove to maintain full generality in our treatment, it does suffer from some limitations.
First, the most glaring limitation is that we only considered the linearized Einstein equations (1.2) in vacuum. Metric reconstruction in the presence of sources is just as interesting (if not more) and has immediate practical implications for the problem of self-force, which bears strong relevance to current and planned gravitational-wave observations. Progress along those lines has been mostly focused on the Schwarzschild case [34, 35], with some notable exceptions [36, 37, 38, 39] beyond the work of Ori [26] and van de Meent and Shah [27].
The non-vacuum case poses at least two new difficulties. First, the Weyl scalars and their associated Hertz potentials no longer obey the same equation: while in the vacuum case, they all solved the homogeneous form (1) of the Teukolsky master equation, in the presence of sources, they must now solve its inhomogeneous version, but with different sources. [^9] Moreover, the reconstruction of a Weyl scalar from the other now involve the source terms, unlike in Sec. 2.6. Though challenging to deal with, this complication is not insuperable. The second difficulty, however, is less easily overcome: the radiation gauges (1.19) and (1.21) cease to work in the presence of sources. More precisely, any metric perturbation satisfying the gauge conditions (1.19) or (1.21) will be singular in some neighborhood of the source [40, 26]. This phenomenon occurs because, when they are combined with the non-vacuum Einstein equations, these gauge conditions place constraints on $T_{\mu\nu}$ that cannot in general be met: even in the simplest case of a stationary point-like source in flat spacetime, the metric perturbation must develop a string-like singularity extending from the source particle [40]. Pound, Merlin, and Barack [41] classified the different types of radiation gauge singularities for a point mass in a general spacetime. They also developed a method for stitching together two gauges that are regular in two different regions to construct a gauge with no string singularities, but at the cost of introducing a jump discontinuity across a surface intersecting the worldline of the particle; additional subtleties also arise when computing any dynamical quantities in this gauge [41]. van de Meent and Shah [27] used this procedure to compute the gravitational waves emitted by a particle on a non-circular equatorial orbit around a Kerr black hole, and van de Meent [37] later extended the analysis to generic Kerr orbits. To our knowledge, an analogous procedure for non-point-like sources has yet to be developed.
@@ 2 Statement of results
After some preliminary definitions in Sec. 2.1, we describe the angular modes $S_{\omega\ell m}^{(s)}(\theta)$ of the Kerr geometry in Sec. 2.2 and then its radial modes $R_{\omega\ell m}^{(s)}(r)$ in Sec. 2.3. Next, in Sec. 2.4, we introduce the corresponding angular constants $D_{\omega\ell m}$, $D_{\omega\ell m}^{\prime}$, and $\mathcal{D}_{\omega\ell m}$, together with radial constants $\mathscr{C}_{\omega\ell m}$, $\mathscr{C}_{\omega\ell m}^{\prime}$, and $\mathcal{C}_{\omega\ell m}$, which appear in the angular and radial Teukolsky–Starobinsky identities that are presented in Sec. 2.5 along with their unseparated analogue. These identities can be used to check the physical relations between the Weyl scalars $\psi_{0}$ and $\psi_{4}$ given in Sec. 2.6. At last, we provide explicit metric reconstruction formulas from the modes of either Weyl scalar in Sec. 2.7 for ingoing radiation gauge and Sec. 2.8 for outgoing radiation gauge. Finally, we discuss some consistency checks in Sec. 2.9, before concluding in Sec. 2.10 with a discussion of the various forms of the Teukolsky–identites and their logical interrelations.
@@ 2.2 Angular modes
As $a\to 0$ and spherical symmetry is restored, the spin-weighted spheroidal harmonics $Z_{\omega\ell m}^{(s)}(\theta,\phi)$ reduce to the standard spin-weighted spherical harmonics, $Y_{\ell m}^{(s)}(\theta,\phi)=N_{\ell m}^{(s)}P_{\ell m}^{(s)}(\theta)e^{im\phi}$, where
$P^{(s)}_{\ell m}(\theta)=\sin^{2\ell}\mathopen{}\mathclose{{}\left(\frac{\theta}{2}}\right)\sum_{k=0}^{\ell-s}(-1)^{k}\binom{\ell-s}{k}\binom{\ell+s}{k+s-m}\cot^{2k+s-m}\mathopen{}\mathclose{{}\left(\frac{\theta}{2}}\right),$ (2.13)
are the “spin-weighted associated Legendre polynomials” and the normalization is [^12] This choice follows from the orthonormality condition (2.12a), that is, $\int_{S^{2}}Y^{(s)}_{\ell m}(\Omega)\overline{Y}^{(s)}_{\ell^{\prime}m^{\prime}}(\Omega)\mathop{}\!\mathrm{d}\Omega=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}$.
$N_{\ell m}^{(s)}=(-1)^{\ell+m-s}\sqrt{\frac{2\ell+1}{4\pi}\frac{(\ell+m)!}{(l+s)!}\frac{(\ell-m)!}{(l-s)!}}.$ (2.14)
In this spinless limit, the eigenvalues $\lambda_{\omega\ell m}^{(s)}$ lose their frequency-dependence and degenerate to

### v2

@@ 1 Introduction and summary
As an aside, we briefly comment on the origin of the two “miracles” uncovered by Teukolsky: 1) the Kerr metric perturbations $h_{\mu\nu}$ are encoded in two gauge-invariant Weyl scalars (1.4), each obeying the same decoupled scalar equation (1), and 2) this equation is completely separable. In fact, Teukolsky showed that the first “miracle” occurs in any algebraically special spacetime of Petrov type D [7]: in this sense, therefore, the reduction of the spin-$s$ field equations to a pair of decoupled scalar equations on $\Psi^{(s)}$ can be viewed as a consequence of algebraic specialness. [^5] In general, this decoupling property only holds for fixed $|s|$. When perturbations of multiple spins are present, their equations may not fully decouple: for example, the Kerr-Newman spacetime is of Petrov type D, but the Teukolsky equations for electromagnetic ($|s|=1$) and gravitational ($|s|=2$) perturbations do not decouple. In Schwarzschild, the second “miracle”—the separability of the spin-$s$ field equations—follows from spherical symmetry. This symmetry is broken in Kerr, but as Carter discovered [8], it is replaced by a “hidden” symmetry in the form of an irreducible rank-2 Killing tensor $K_{\mu\nu}$, whose existence implies the separability of the scalar wave equation. Indeed, the Killing equation for $K_{\mu\nu}$ implies that the differential operator $\mathcal{K}=\nabla^{\mu}K_{\mu\nu}\nabla^{\nu}$ commutes with the Laplacian $\nabla^{2}=\nabla^{\mu}g_{\mu\nu}\nabla^{\nu}$, and both commute with the action of the Killing vectors $\mathop{}\!\partial_{t}$ and $\mathop{}\!\partial_{\phi}$ generating the isometries of Kerr (stationarity and axisymmetry, respectively); since these four differential operators all commute, they are simultaneously diagonalizable, and their joint eigenvectors are precisely the separated eigenmodes $\Psi_{\omega\ell m}^{(0)}=e^{-i\omega t+im\phi}R_{\omega\ell m}^{(0)}(r)S_{\omega\ell m}^{(0)}(\theta)$; [^6] The eigenvalues corresponding to $i\mathop{}\!\partial_{t}$, $i\mathop{}\!\partial_{\phi}$, $\nabla^{2}$, and $\mathcal{K}$ are respectively the energy $\omega$, azimuthal angular momentum $m$, invariant mass $\mu^{2}$, and separation constant $\lambda_{\omega\ell m}^{(0)}$ of the mode solution to $\mathopen{}\mathclose{{}\left(\nabla^{2}-\mu^{2}}\right)\Psi_{\omega\ell m}^{(0)}=0$. Likewise, the geodesic equation (which is the eikonal limit of the wave equation) is completely integrable due to the existence of four conserved quantities associated with these operators (e.g., $\mathcal{K}_{\mu\nu}p^{\mu}p^{\nu}$ is the Carter constant). see, e.g.,[9] for a more detailed review. More generally, the separability of the spin-$s$ field equations can be derived from the existence of commuting differential operators, all of which are “hidden” in the sense that they are associated with symmetries of higher rank, rather than being geometrically realized as isometries. [^7] The symmetry operators that separate the spin-$s$ equations in Kerr are derived by Carter and McLenaghan [10] for $|s|=\frac{1}{2}$ and by Torres del Castillo [11] for $|s|=1$. Aksteiner and Bäckdahl [12] tackled the case $|s|=2$. While satisfying, symmetry-based arguments do not fully explain these “miracles”: for one, the existence of “hidden” symmetries is sufficient but not necessary for separability. [^8] The C-metric describing a pair of accelerating black holes [13] provides a counter-example: it is algebraically special of Petrov type D and therefore admits a conformal Killing tensor, but it does not possess an exact one [14]. We are not aware of any method to derive separability from conformal symmetries and their differential operators. Dudley and Finley [15] proved that all spacetimes in the Plebański-Demiański class [16], which includes all vacuum type D solutions, share these two properties (decoupling and separability).
@@
Although knowledge of the Weyl scalars $\psi_{0}$ and/or $\psi_{4}$ is adequate for many purposes (such as the computation of energy and angular momentum fluxes), the full metric perturbation $h_{\mu\nu}$ is often required to determine other aspects of the perturbed spacetime geometry (such as the induced metric on the horizon), or to push perturbation theory beyond first order [18, 19]. It is therefore useful to reconstruct the specific metric perturbation $h_{\mu\nu}$ associated with a given Weyl scalar. Various methods for carrying out this metric reconstruction have been developed, but the explicit form of the metric perturbation $h_{\mu\nu}$ that corresponds to a single mode of $\psi_{0}$ or $\psi_{4}$ has—to our knowledge—never been published. The main purpose of this paper is to close this gap in the literature. Before describing our results, we briefly summarize preexisting work.
@@
$\text{ORG}:\qquad n^{\mu}h_{\mu\nu}^{\rm ORG}=0,\qquad g^{\mu\nu}h_{\mu\nu}^{\rm ORG}=0.$ (1.21)
@@ 1.1 Mode inversion and comparison with previous work
To carry out the metric reconstruction procedure explicitly, one must invert the relations (1.22) and (1.23) to obtain the Hertz potentials $\Psi_{\rm H}$ and $\Psi_{\rm H}^{\prime}$ in terms of the Weyl scalars $\psi_{0}$ and $\psi_{4}$. Although this inversion is difficult—or perhaps even impossible—to perform in full generality, it is straightforward (if messy) to invert these equations at the mode level, that is, for the case where $\Psi^{(+2)}=\psi_{0}$ or $\Psi^{(-2)}=\zeta^{4}\psi_{4}$ consists of a single mode (1.9b). For real frequencies $\omega$, the results of this mode inversion have appeared previously in the literature, though scattered across multiple papers [26, 27, 28, 29], as we review below. Here, we tackle the general case of arbitrary (and possibly complex) frequency $\omega$, and completely invert all of the relations (1.22) and (1.23).
@@
In their excellent review, Pound and Wardell [28] inverted precisely the other two relations. Their Eq. (106) inverts our Eq. (1.23a) for $\Psi_{\rm H}^{\prime}$ (in their notation, $\psi^{\rm ORG}$) in terms of $\Psi^{(+2)}=\psi_{0}$, which they denote $\psi_{2}$—see their Eq. (81). Likewise, their Eq. (107) inverts our Eq. (1.22b) for $\Psi_{\rm H}$ (in their notation, $\psi^{\rm IRG}$) in terms of $\Psi^{(-2)}=\zeta^{4}\psi_{4}$, which they label $\psi_{-2}$—see their Eq. (82). Together, the three papers [26, 27, 28] completely invert all four of the relations (1.22) and (1.23), but only for the case of real frequency $\omega$, which can be a serious limitation for many applications. An earlier work [29] also inverted one of these relations for complex $\omega$ (see App. C therein).
Here, we invert all four of these relations for complex frequency $\omega$. As a technical aside, there is another redundancy inherent in mode inversion that can simplify the problem. The key observation is that, even if one wishes to use both $\psi_{0}$ and $\psi_{4}$, and work in both IRG and ORG, it is still enough to invert only two of the relations (1.22) and (1.23): that is, it suffices to invert Eqs. (1.22b) and (1.23a) only, as Pound and Wardell did, or else Eqs. (1.22a) and (1.23b) only.
@@
The second step was completely worked out for real frequency $\omega$ [26, 27, 28], and in one case also for complex $\omega$ [29]. Here, we extend mode inversion with complex frequency $\omega$ to all cases. We also derive direct relations between modes of $\psi_{0}$ and $\psi_{4}$, which are to our knowledge novel.
@@ 1.2 Structure of the paper
Lastly, we summarize the theory of the confluent Heun function, and carefully derive the quantization conditions for its eigenvalues, in App. E. We also discuss the orthogonality of the confluent Heun functions and derive an analytic formula for their normalization, following the work of Becker [34] on the full Heun equation. Finally, we apply these results to the angular ODE (1.11) in App. F and to the radial ODE (1.10) in App. G. Beyond this paper, we also provide Mathematica notebooks containing our formulas and their checks on this Github.
@@ 1.3 Present limitations and future directions
While we strove to maintain full generality in our treatment, it does suffer from some limitations. First, the most glaring limitation is that we only considered the linearized Einstein equations (1.2) in vacuum. Metric reconstruction in the presence of sources is just as interesting (if not more) and has immediate practical implications for the problem of self-force, which bears strong relevance to current and planned gravitational-wave observations. Progress along those lines has been mostly focused on the Schwarzschild case [35, 36], with a few notable exceptions [37, 38, 39, 40] beyond the work of Ori [26] and van de Meent and Shah [27].
The non-vacuum case poses at least two new difficulties. First, the Weyl scalars and their associated Hertz potentials no longer obey the same equation: while in the vacuum case, they all solved the homogeneous form (1) of the Teukolsky master equation, in the presence of sources, they must now solve its inhomogeneous version, but with different sources. [^12] Moreover, the reconstruction of a Weyl scalar from the other now involves the source terms, unlike in Sec. 2.6. Though challenging to deal with, this complication is not insuperable. The second difficulty, however, is less easily overcome: the radiation gauges (1.19) and (1.21) cease to work in the presence of sources. More precisely, any metric perturbation satisfying the gauge conditions (1.19) or (1.21) will be singular in some neighborhood of the source [41, 26]. This phenomenon occurs because, when they are combined with the non-vacuum Einstein equations, these gauge conditions place constraints on $T_{\mu\nu}$ that cannot in general be met: even in the simplest case of a stationary point-like source in flat spacetime, the metric perturbation must develop a string-like singularity extending from the source particle [41]. Pound, Merlin, and Barack [42] classified the different types of radiation gauge singularities for a point mass in a general spacetime. They also developed a method for stitching together two gauges that are regular in two different regions to construct a gauge with no string singularities, but at the cost of introducing a jump discontinuity across a surface intersecting the worldline of the particle. van de Meent and Shah [27] used this procedure to compute the gravitational waves emitted by a particle on a non-circular equatorial orbit around a Kerr black hole, and van de Meent [38] later extended the analysis to generic Kerr orbits. To our knowledge, an analogous procedure for non-point-like sources has yet to be developed.
@@ 2 Statement of results
After some preliminary definitions in Sec. 2.1, we describe the angular modes $S_{\omega\ell m}^{(s)}(\theta)$ of the Kerr geometry in Sec. 2.2 and then its radial modes $R_{\omega\ell m}^{(s)}(r)$ in Sec. 2.3. Next, in Sec. 2.4, we introduce the corresponding angular constants $D_{\omega\ell m}$, $D_{\omega\ell m}^{\prime}$, and $\mathcal{D}_{\omega\ell m}$, together with radial constants $\mathscr{C}_{\omega\ell m}$, $\mathscr{C}_{\omega\ell m}^{\prime}$, and $\mathcal{C}_{\omega\ell m}$, which appear in the angular and radial Teukolsky–Starobinsky identities that are presented in Sec. 2.5 along with their unseparated analogue. These identities can be used to check the physical relations between the Weyl scalars $\psi_{0}$ and $\psi_{4}$ given in Sec. 2.6. At last, we provide explicit metric reconstruction formulas from the modes of either Weyl scalar in Sec. 2.7 for ingoing radiation gauge and Sec. 2.8 for outgoing radiation gauge. Finally, we discuss some consistency checks in Sec. 2.9, before concluding in Sec. 2.10 with a discussion of the various forms of the Teukolsky–Starobinsky identities and their logical interrelations.
@@ 2.2 Angular modes
As $a\to 0$ and spherical symmetry is restored, the spin-weighted spheroidal harmonics $Z_{\omega\ell m}^{(s)}(\theta,\phi)$ reduce (up to signs) to their standard spherical counterparts $Y_{\ell m}^{(s)}(\theta,\phi)=P_{\ell m}^{(s)}(\cos{\theta})e^{im\phi}$, where
$P_{\ell m}^{(s)}(\cos{\theta})=N_{\ell m}^{(s)}\sin^{2\ell}\mathopen{}\mathclose{{}\left(\frac{\theta}{2}}\right)\sum_{k=0}^{\ell-s}(-1)^{k}\binom{\ell-s}{k}\binom{\ell+s}{k+s-m}\cot^{2k+s-m}\mathopen{}\mathclose{{}\left(\frac{\theta}{2}}\right),$ (2.13)
are “spin-weighted associated Legendre polynomials” with normalization [^15] This choice follows from the orthonormality condition (2.12a), that is, $\int_{S^{2}}Y^{(s)}_{\ell m}(\Omega)\overline{Y}^{(s)}_{\ell^{\prime}m^{\prime}}(\Omega)\mathop{}\!\mathrm{d}\Omega=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}$.
$N_{\ell m}^{(s)}=(-1)^{\ell+m-s}\sqrt{\frac{2\ell+1}{4\pi}\frac{(\ell+m)!}{(\ell+s)!}\frac{(\ell-m)!}{(\ell-s)!}}.$ (2.14)
In the limit $a\to 0$, the eigenvalues $\lambda_{\omega\ell m}^{(s)}$ lose their frequency-dependence and degenerate to